help!Two particles, one with charge -2.09 uC and one with charge 6.91 C, are 6.59 centimeters apart. What is the magnitude of the force that one particle exerts on the other? Number Two new particles, which have identical positive charge q3, are placed the same 6.59 centimeters apart, and the force between them is measured to be the same as that between the original particles. What is q3? Number uC
Answer #1

1) By using Coulombs law,

Force,F = K*q1*q2/r^2

= 9*10^9 Nm^2/C^2 *(-2.09 C)*(6.91 C)/(0.0659) m^2

= 1.97*10^12 N